Circular No. 7374

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000N IN MCG –02-34-054 Y. Sato and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery of an apparent supernova (mag ∼ 17.3) on an unfiltered image taken on Mar. 4.4 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2000N is located at α = 13^{h}27^{m}56^{s}.19, δ = −13^{o}24′48′′.3 (equinox 2000.0), which is 9′′.2 east and 34′′.1 north of the nucleus of MCG –02-34-054. SN 2000N was confirmed at about the same magnitude on an unfiltered image taken by T. Puckett with the Puckett Observatory 0.60-m automated supernova patrol telescope on Mar. 5.2. An unfiltered KAIT image taken on 1999 Dec. 25.5 showed nothing at the position of the new object (limiting mag ∼ 19.0).

SUPERNOVA 2000M IN NGC 6389 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000M (cf. IAUC 7373), taken by M. Calkins on Mar. 2.5 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), reveals it to be a type-II supernova after maximum light. The spectrum exhibits a slightly red continuum with well-developed lines, including Fe II, Na I, and P-Cyg Hα. Emission lines from the host galaxy yield a recession velocity of 3160 ± 20 km/s. The supernova expansion velocity, measured using the Hα line profile, is 5600 km/s.

E. Cappellaro, M. Turatto, S. Benetti, and A. Pastorello, Padua, observed SN 2000M with the European Southern Observatory 3.6-m reflector (+ EF2) on Mar. 3.4 UT. Preliminary reduction of the spectrum (range 350–750 nm, resolution 15 nm) shows strong lines with P-Cyg profiles over a red continuum. The spectrum resembles that of a type-II supernova, ∼ 1 month past maximum, with strong lines of H I, Na I D, Ba II, Sc II, and Fe I. The photospheric expansion velocity, determined from the minimum of the Hα absorption, is ∼ 6000 km/s. The red continuum seems intrinsic and also due to foreground extinction in the Galaxy and in the parent galaxy. In fact, a relatively strong absorption feature is visible at 595 nm, corresponding to the Na I D doublet at the parent-galaxy redshift.

2000 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green